a brass rod is 0.560 in long at 23.0 oc what is the length of the rod if it is heated to 57.0 oc?


1
Expert's answer
2021-04-30T11:23:44-0400
"l=l_o(1+\alpha \Delta T)\\l=0.56(1+19\cdot10^{-6}(57-23))=0.56036\ in"